# VectorLiteRAG: Latency-Aware and Fine-Grained Resource Partitioning for Efficient RAG

### v2

@@ V-A Profiling-based Performance Modeling
Since GPU resources are limited, accurately modeling system performance is critical for determining the optimal index partitioning point. To construct these models, VectorLiteRAG profiles latency and access statistics using calibration queries from a training set. Specifically, it collects: (1) latency breakdown of CPU-based vector search and (2) cluster access frequency distributions. Additionally, the standalone throughput of the LLM is measured to guide partitioning decisions under joint CPU-GPU execution.
@@
As described in Section II-B, IVF index search latency is dominated by two components: coarse quantization (CQ) and LUT operations. We profile both stages across varying batch sizes and construct independent models for each. However, in our design, only the LUT stage is considered for GPU offloading, for two main reasons:
@@
First, CQ is itself a similarity search over the quantizer (centroid) vectors, which is often implemented using memory-intensive graph-based structures such as HNSW. Offloading CQ to GPU would require additional memory for the graph and complicate memory management. Second, if CQ were distributed across GPU shards, the resulting search path would involve repeated device transitions: CPU $\rightarrow$ GPU (quantization) $\rightarrow$ CPU (merge and routing) $\rightarrow$ GPU (search) $\rightarrow$ CPU (final merge). This induces costly inter-device communication and synchronization overheads. To avoid these complications, we retain CQ on the CPU and focus GPU optimization on LUT operations.
@@
When hot clusters are cached, overall search time reduces accordingly. In particular, since LUT operations offloaded to the GPU they are fully overlapped with the remaining CPU-side processing. As a result, we model the latency of the hybrid partitioned index as:
@@
where $\eta$ denotes the hit rate, more precisely the minimum hit rate among all queries in the batch.
V-B Tail Query Hit Rate Estimation
As discussed in Section III-C, caching hot clusters leads to varying hit rates across queries. Since LUT workload is proportional to the number of clusters processed on the CPU ($1-\eta$), this variance directly translates to latency differences.
To maximize throughput, vector search is typicall y executed in batches. However, batch-level search latency is dictated by the slowest, tail query. Therefore, modeling the minimum hit rate within a batch is critical for accurate performance estimation.
We model the distribution of per-query hit rates using a Beta distribution $f(x)$, which is widely used in Bayesian statistics to model random variables constrained between 0 and 1. Given a batch of size $b$, the expected minimum hit rate $\eta_{\min}$ (i.e., the first-order statistic) is computed as:
@@ V-B Tail Query Hit Rate Estimation
where $F(x)$ is the cumulative distribution function (CDF) of $f(x)$.
The mean hit rate $\bar{\eta}$ is easily derived from the query-cluster access frequency profile: the cumulative access percentage of cached clusters. Estimating the variance is more challenging, as it would require re-running queries through the quantizer and counting individual hits after masking hot clusters, a process that is both computationally expensive and incompatible with iterative partitioning algorithm.
Instead, we approximate the hit rate variance as a function of the mean. We observe that hit/miss heterogeneity is maximized when $\bar{\eta}=0.5$, and becomes more uniform as $\bar{\eta}\to 0$ or $\bar{\eta}\to 1$. This mirrors the variance behavior of the Beta distribution, where:
$\text{Var}(X)\propto\bar{\eta}(1-\bar{\eta})$
Thus, by empirically profiling the variance at $\bar{\eta}=0.5$, denoted $\sigma_{\max}^{2}$, we can approximate the variance at arbitrary $\bar{\eta}$ as:
@@
Figure 8 (right) validates the approximation. This allows instantiating a Beta distribution with inferred mean and variance for any cache coverage configuration.
@@
V-C Latency-Bounded Partitioning Algorithm
In proposed hybrid CPU-GPU RAG pipeline, LLM throughput is inversely related to vector index cache coverage, due to GPU memory contention between KV cache and index storage. To find the optimal balance, we propose an iterative algorithm that searches for the optimal index partitioning point.
Algorithm 1 outlines the proposed latency-bounded partitioning algorithm. It takes the following inputs: the latency target, the baseline KV cache memory footprint when no vector index is loaded, and the peak standalone LLM throughput. The goal is to determine the largest feasible GPU index cache coverage (partitioning point $\rho$) that satisfies the latency constraint.
We begin by computing the latency bound for the hybrid vector search stage. To account for queuing delay, we conservatively consider the worst-case scenario: a request arriving just after the previous batch has started processing. Under steady-state load with uniformly arriving requests, this tail query experiences the full batch latency $W(b)$ as queuing delay. Hence, to ensure total response time remains within the latency budget, we bound the hybrid search latency as:
$\tau_{s}=\text{SLO}_{\text{search}}-W(b)$ (3)
To simplify analysis and avoid circular dependency (since $W(b)$ depends on $\tau_{s}$), we approximate this using a queuing factor $\epsilon$, leading to:
@@ V-C Latency-Bounded Partitioning Algorithm
In our setting, we conservatively set $\epsilon=1$, corresponding to the worst-case queuing delay equal to one batch latency.
Next, we perform binary search over possible values of $\rho$ using the modeled latency and hit rate behavior.
Search iteration. For each candidate $\rho$, we first estimate the reduced LLM throughput due to decreased KV cache space. While this interpolation is coarse, it gives a conservative lower bound because the throughput–cache curve is generally convex. We then invoke the InferPartition function. It computes the expected batch size as $B=\mu\cdot\tau_{s}$, where $\mu$ is the current throughput bound. Since batch size $B$ must be an integer, we consider two rounding strategies:
@@
6:  $\rho\leftarrow\textsc{InferPartition}(t_{s},\mu_{\text{LLM}})$
@@
17:  $T^{\text{CPU}}_{\text{search}}(B),T^{\text{CPU}}_{\text{LUT}}(B)\leftarrow\textsc{PerfModel}(B)$
@@
19:  $\rho_{1}\leftarrow\textsc{Hitrate2Coverage}(\eta_{1},B)$
@@
21:  $T^{\text{CPU}}_{\text{search}}(B),T^{\text{CPU}}_{\text{LUT}}(B)\leftarrow\textsc{PerfModel}(B)$
@@
23:  $\rho_{2}\leftarrow\textsc{Hitrate2Coverage}(\eta_{2},B)$
@@
We then select the smaller of $\rho_{1}$ and $\rho_{2}$ to minimize GPU memory usage. This value is used to update the binary search interval.
Convergence. If the new partitioning point $\rho$ increases, the resulting decrease in throughput leads to a smaller batch size in the next iteration, pulling $\rho$ back down. Conversely, if $\rho$ shrinks, the throughput bound increases, allowing for more cache coverage. This feedback loop ensures convergence of the binary search within a small number of iterations.
In practice, the algorithm converges in under one minute and provides an optimal GPU–CPU partitioning point $\rho$ that satisfies the latency target while preserving maximum LLM throughput.
V-D Index Splitter
Once the partitioning point $\rho$ is determined, it is passed to the final stage of index construction: the index splitter. The splitter first identifies the hot clusters based on the access profile and the target cache coverage $\rho$. These hot clusters are then sorted by size and distributed to GPU shards in a round-robin fashion to balance memory usage across sub-indexes.
VI Distributed VectorLiteRAG Pipeline
The right side of Figure 7 illustrates the runtime architecture of VectorLiteRAG. Similar to other IVF-based indexes, the pipeline begins with coarse quantization to identify candidate clusters. From this point, however, VectorLiteRAG introduces a customized retrieval pipeline tailored for hybrid CPU-GPU execution. We now describe each component in detail.
VI-1 Router
To support efficient retrieval on a distributed multi-GPU system, VectorLiteRAG implements a custom router rather than relying on Faiss’s built-in IndexIVFShards. The default implementation in Faiss is suboptimal in constrained environments for two main reasons:
(1) IndexIVFShards partitions the index uniformly by vector or cluster ID, ignoring access frequency. While, convenient for implementation, it retains centroid metadata even for clusters that are not locally resident, causing unnecessary memory overhead, especially problematic when the number of clusters is large.
@@ VI-1 Router
(2) During search, each sub-index is instructed to probe the same number of clusters, even if many of them are not resident on that shard. Although certain probes are ultimately skipped at runtime, the batched execution of cluster scanning kernels still launches GPU thread blocks for them. These launches consume scheduling bandwidth and shared memory resources, regardless of whether the actual computation is needed. Since shared memory usage increases with nprobe, this results in inefficient kernel launches and exacerbates resource contention, especially in large-scale vector databases.
@@
To address these issues, VectorLiteRAG uses the mapping tables generated during index splitting to compactly route each query to the appropriate GPU shards and prune irrelevant probes. As a result, the effective nprobe per shard is significantly reduced, lowering both memory and kernel scheduling overhead. At runtime, only GPU workers holding relevant clusters receive and execute the search request, while the remaining portion of the search is handled by the CPU. This hybrid execution minimizes contention and ensures efficient utilization of GPU memory and compute.
VI-2 Dynamic Dispatcher
@@
Because hit rates vary across queries, the effective nprobe differs even within a batch. As batch size increases, the minimum hit rate tends to decrease, elongating the search latency for the entire batch. To mitigate this issue, VectorLiteRAG employs a dynamic dispatcher that accelerates early query completion.
When search is initiated, a separate dispatcher thread is launched. Each GPU worker sets a completion flag once its assigned clusters are scanned. After all GPU flags are set, the dispatcher begins polling for queries that have completed their full search.
@@
To facilitate timely query promotion, a callback mechanism connects the CPU search loop and the dispatcher. The CPU processes clusters one-by-one, grouped by related queries. At the end of each iteration, the current scan count is compared with the expected nprobe for each query. When all assigned clusters for a query are scanned, the callback is invoked, and the query and its results are inserted into a thread-safe queue.
@@
The dispatcher polls this queue at short intervals. Once a completed query is available, it merges the CPU and GPU results, re-ranks them to obtain the final top-$k$ vectors, and forwards the result to the downstream document retriever. This proactive execution reduces head-of-line blocking within batches and improves end-to-end latency, particularly for high-hit-rate queries. It also enhances batching continuity by enabling smoother transitions between retrieval and generation stages, which already employs continous batching schemes.
VII Methodology
VII-A Experiment Setup
@@
Datasets and Models. We use two datasets: Wiki-All and ORCAS. IVF index was built following the configuration guide presented in the Faiss library. Wiki-All [35] contains 88M 768-dimensional vectors derived from Wikitext (Kaggle) [26] and Cohere Wikipedia embeddings, with a compressed IVF index occupying 18GB.
Additionally, we construct two indexes from chunked Wikipedia documents using the Stella [39] embedding model of dimensions 1024 and 2048, with queries sampled from the Microsoft ORCAS dataset [4]. ORCAS consists of real Bing queries and retains duplicates to capture realistic query distributions. The ORCAS 1K and ORCAS 2K indexes require 40GB and 80GB of memory, respectively.
Our retrieval pipeline builds on Faiss v1.9.0 [5, 16], with internal extensions for flexible nprobe settings and dispatcher callbacks. The system, including the profiler and latency-aware scheduler, is implemented in Python.
For generation, we use Llama3-8B, Qwen3-32B, and Llama3-70B [6, 38], served via vLLM v0.9.1 [17]. The retriever and LLM servers run as separate subprocesses, while the main process manages request generation and document fetching, enabling seamless integration of the end-to-end RAG pipeline.
To evaluate system performance, we sample queries from a dedicated test set that is disjoint from the profiling set. The request arrival process follows a Poisson distribution, a commonly adopted modeling choice in prior work [17, 42, 29]. For each query, the top-25 documents are retrieved, and a 1024-token input is constructed and passed to the LLM, which then generates a 256-token output, following the setup in [33]. The initial nprobe is set to 2048, which is sufficient to achieve a retrieval quality of 0.91 Normalized Discounted Cumulative Gain (NDCG) [37] at 50.
@@
TABLE I: Latency requirement of each stage used in the evaluation
Figure 9: TTFT SLO attainment of RAG pipeline under increasing arrival rates across different LLMs (columns) and datasets (rows). Our work (vLiteRAG) achieves higher SLO attainment across all regimes compared to baselines, especially under high-load conditions and larger vector index.
@@
System Configuration. Experiments were conducted on two types of nodes, each equipped with eight NVIDIA GPUs. The L40S node includes L40S GPUs with 48GB GDDR memory and dual Xeon 6426Y CPUs. The H100 node uses H100 GPUs with 80GB HBM and Xeon Platinum 8462Y CPUs. We used the L40S node for smaller models (Llama3-8B), while larger models requiring model parallelism (Qwen3-32B, Llama3-70B) were run on the H100 node for maximum throughput.
@@
Baseline Configurations. We compare VectorLiteRAG against several key baselines. Since VectorLiteRAG builds on FAISS, we use vanilla FAISS-CPU IVF FastScan (CPU-Only), FAISS-GPU IVF on a dedicated GPU (DED-GPU), and a sharded FAISS-GPU IVF index distributed across all GPUs (ALL-GPU). To further demonstrate the strength of our approach, we also compare against HeteRAG [10] in section VIII-D, which also uses a skew-aware caching strategy.
VIII Evaluations
VIII-A Performance Model and Hit Rate Estimator

### v3

@@ IV-A1 Profiling-based Performance Modeling
Since GPU resources are limited, accurately modeling performance is critical for determining the optimal index partitioning point. To construct these models, VectorLiteRAG profiles latency and access statistics using calibration queries from a training set. Specifically, it collects: (1) latency breakdown of CPU-based vector search and (2) cluster access frequency distributions. Additionally, throughput of the bare LLM is measured to guide partitioning decisions under joint CPU-GPU execution.
@@
As described in Section II-B, IVF index search latency is dominated by two components: coarse quantization (CQ) and LUT operations. We profile both stages across varying batch sizes and construct independent models for each. However, in our design, only the LUT stage, which corresponds to the individual distance computation and scanning step, is considered for GPU offloading for two main reasons:
@@
First, CQ is a similarity search over the quantizer (centroid) vectors, which is often implemented using memory-intensive graph-based structures such as HNSW. Offloading CQ to GPU would require additional memory for the graph and complicate memory management. Second, if CQ were distributed across GPU shards, the resulting search path would involve repeated device transitions: CPU $\rightarrow$ GPU (quantization) $\rightarrow$ CPU (merge and routing) $\rightarrow$ GPU (search) $\rightarrow$ CPU (final merge). This induces costly inter-device communication and synchronization overheads. Moreover, our objective is to ensure stable performance within the latency budgets rather than to minimize absolute latency. Thus, for our purpose, we retain CQ on the CPU and use GPUs for distance computations, as this offers performance benefits while simplifying the optimization space.
@@
When hot clusters are cached, the overall search time reduces accordingly. LUT operations offloaded to GPUs are fully hidden under CPU’s execution, and the CPU processing time decreases in proportion to the number of hits. As a result, we model the latency of the hybrid partitioned index as:
@@
where $\eta$ denotes the hit rate, in particular the minimum hit rate among all queries in the batch.
IV-A2 Tail Query Hit Rate Estimation
As discussed in Section III-C, caching hot clusters leads to varying hit rates across queries. Because, CPU side LUT workload is proportional to the miss rate ($1-\eta$), this variance directly translates into differences in search latency. Moreover, since vector search is typically executed in batches to maximize throughput, the completion time of the entire batch is dictated by the slowest query, one with the fewest hits. Therefore, modeling the minimum hit rate within a batch is critical for accurate performance estimation.
We model the distribution of per-query hit rates using a Beta distribution $f(x)$, which is widely used in Bayesian statistics for variables constrained to the $[0,1]$ range. For a batch of size $b$, the expected minimum hit rate $\eta_{\min}$, i.e., the first-order statistic, is computed as:
@@ IV-A2 Tail Query Hit Rate Estimation
where $F(x)$ is the cumulative distribution function of $f(x)$.
The mean hit rate $\bar{\eta}$ can be obtained directly from the query–cluster access profile, which reflects the cumulative fraction of accesses covered by the cached clusters. Estimating the variance is more challenging, as it would require re-running queries through the quantizer and counting individual hits after masking hot clusters, a process that is both computationally expensive and incompatible with iterative partitioning algorithm.
Instead, we approximate the hit rate variance as a function of the mean. We observe that hit/miss variance peaks when $\bar{\eta}=0.5$, and becomes more uniform as $\bar{\eta}\to 0$ or $\bar{\eta}\to 1$. This mirrors the variance behavior of the Beta distribution; $\text{Var}(X)\propto\bar{\eta}(1-\bar{\eta})$. Thus, by empirically profiling the variance at $\bar{\eta}=0.5$, denoted $\sigma_{\max}^{2}$, we can approximate the variance at arbitrary $\bar{\eta}$ as:
@@
Figure 8 (right) validates the approximation. This allows instantiating a Beta distribution $f(x)$ with inferred mean and variance for any cache coverage configuration.
@@
IV-A3 Latency-Bounded Partitioning Algorithm
In the hybrid RAG pipeline, LLM throughput decreases as more GPU memory is allocated to the vector index, due to contention between KV cache and index storage. To balance these competing demands, we introduce an iterative algorithm that determines an index partitioning point satisfying the latency constraint.
Algorithm 1 outlines the proposed latency-bounded partitioning algorithm. It takes the following inputs: the latency target, the baseline KV cache memory footprint when no vector index is loaded, and the peak bare LLM throughput. The goal is to find the largest feasible cache coverage for the GPU index (partitioning point $\rho$) that satisfies SLO constraint.
We first compute the latency bound for the hybrid vector search stage. To account for queuing delay, the analysis considers a worst-case scenario in which a request arrives immediately after the previous batch begins processing. Under steady-state load with uniformly arriving requests, this tail query experiences full batch latency $W(b)$ as queuing delay. To maintain the total response time within the latency budget, the search latency must satisfy $\tau_{s}\leq\text{SLO}_{\text{search}}-W(b)$.
To avoid circular dependency (as $W(b)$ depends on $\tau_{s}$), we approximate this term using a queuing factor $\epsilon$, leading to:
@@ IV-A3 Latency-Bounded Partitioning Algorithm
In our setting, we set $\epsilon=1$, as it represents the worst case where the queuing delay equals one batch latency. This choice is empirically supported from the CPU-only baseline, where $\epsilon$ ranged between 0.9 and 1.0.
Search iteration. The algorithm then performs a binary search over possible values of $\rho$ using the modeled latency and hit rate behavior. For each candidate $\rho$, the reduced LLM throughput is estimated based on the corresponding decrease in KV cache capacity. Although this interpolation is coarse, it provides a conservative lower bound because the throughput–cache curve is generally convex. The InferPartition function is subsequently invoked to compute the expected batch size, given by $B=\mu\cdot\tau_{s}$, where $\mu$ is the current throughput bound. Since batch size $B$ must be an integer, two rounding strategies are considered:
@@
At the end of the iteration, the smaller of $\rho_{1}$ and $\rho_{2}$ is selected, as it requires less GPU memory. This value is used to update the binary search interval.
Convergence. If the newly computed partitioning point $\rho$ increases, the resulting drop in throughput leads to a smaller batch size in the next iteration, which in turn drives $\rho$ back down. Conversely, if $\rho$ shrinks, the throughput bound increases, allowing for more cache coverage. This feedback loop ensures convergence of the algorithm within a limited number of iterations. In practice, convergence takes less than one minute as shown in Figure 9.
IV-A4 Index Splitter
Once the partitioning point $\rho$ is determined, it is passed to the final stage of index construction, which is the index splitter. The splitter first identifies the hot clusters based on the access profile and the target cache coverage $\rho$. These hot clusters are then sorted by size and distributed to GPU shards in a round-robin fashion to balance memory usage across sub-indexes.
IV-B Distributed VectorLiteRAG Pipeline
The right side of Figure 7 illustrates the runtime architecture of VectorLiteRAG. At initialization, memory is allocated sequentially for the index and then for the LLM to prevent memory interference between the vector search and LLM engines. The two components operate through different processes and thus use separate GPU streams for concurrency.
Similar to other IVF-based indexes, the pipeline begins with coarse quantization to identify candidate clusters. However, from this point on, VectorLiteRAG introduces a customized retrieval pipeline tailored for hybrid CPU-GPU execution. We now describe each component in detail.
@@ IV-B Distributed VectorLiteRAG Pipeline
6:  $\rho\leftarrow\text{{InferPartition}}(t_{s},\mu_{\text{LLM}})$
@@
17:  $T^{\text{CPU}}_{\text{search}}(B),T^{\text{CPU}}_{\text{LUT}}(B)\leftarrow\text{{PerfModel}}(B)$
@@
19:  $\rho_{1}\leftarrow\text{{Hitrate2Coverage}}(\eta_{1},B)$
@@
21:  $T^{\text{CPU}}_{\text{search}}(B),T^{\text{CPU}}_{\text{LUT}}(B)\leftarrow\text{{PerfModel}}(B)$
@@
23:  $\rho_{2}\leftarrow\text{{Hitrate2Coverage}}(\eta_{2},B)$
@@
IV-B1 Router
@@
To support efficient vector retrieval on a distributed multi-GPU system, VectorLiteRAG implements a custom routing mechanism rather than relying on Faiss’s built-in IndexIVFShards. The default implementation in Faiss is suboptimal in constrained environments for two main reasons. (1) IndexIVFShards partitions the index uniformly by vector or cluster ID, ignoring access frequency. While, convenient for implementation, it retains centroid metadata even for clusters that are not locally resident, causing unnecessary memory overhead, especially problematic when the number of clusters is large. (2) During search, each sub-index is instructed to probe the same number of clusters, even if many of them are not resident on that shard. Although certain probes are ultimately skipped at runtime, the batched execution of cluster scanning kernels still launches GPU thread blocks for them. These launches consume scheduling bandwidth and shared memory resources, regardless of whether the actual computation is needed. Since shared memory usage increases with nprobe, this results in inefficient kernel launches and exacerbates resource contention, especially in large-scale vector databases.
@@
To address these issues, VectorLiteRAG uses the mapping tables generated during index splitting to route each query to the appropriate GPU shards and prune irrelevant probes, thereby accounting for the device-level variance. This substantially reduces the effective nprobe per shard, lowering both memory pressure and kernel scheduling overhead. At runtime, only GPU workers holding relevant clusters receive and execute the search request, while the remaining portion of the search is handled by the CPU. This hybrid execution minimizes contention and enables more efficient use of GPU memory and compute resources.
IV-B2 Dynamic Dispatcher
@@
Because hit rates vary across queries, the effective nprobe differs even within a batch. As batch size increases, the minimum hit rate tends to decrease, increasing the search latency for the entire batch. To mitigate this issue, VectorLiteRAG employs a dynamic dispatcher that accelerates early query completion.
@@
When search is initiated, a separate dispatcher thread is launched. Each GPU worker sets a completion flag once its assigned clusters are scanned. After all GPU flags are set, the dispatcher begins polling for queries that have completed their full search. To facilitate timely query promotion, a callback mechanism connects the CPU search loop and the dispatcher, as CPU processes clusters one-by-one, grouped by related queries. At the end of each iteration, the current scan count is compared with the expected nprobe for each query. When all assigned clusters for a query are scanned, the callback is invoked, and the query and its results are inserted into a thread-safe queue.
@@
The dispatcher polls this queue at short intervals. Once a completed query is available, it merges the CPU and GPU results, re-ranks them to obtain the final top-$k$ vectors, and forwards the result to the downstream document retriever. This proactive execution reduces head-of-line blocking within batches and improves end-to-end latency, particularly for high-hit-rate queries. It also enhances batching continuity by enabling smoother transitions between retrieval and generation stages, which already employs continuous batching schemes.
IV-B3 Adaptive Runtime Index Update
Our model is built upon the distributional characteristics of queries aggregated across batches. While correlations among queries may temporarily shift access patterns, they primarily reduce the number of statistically independent samples rather than altering the overall distributional trend. Nevertheless, temporal bias can arise in practice, and to mitigate potential performance degradation caused by such drift, VectorLiteRAG employs an adaptive re-profiling and update process.
VectorLiteRAG can swiftly react to shifts in query distribution without interrupting service. During runtime, the router monitors (1) average hit rates and (2) per-cluster access frequencies. For every few minutes or after a few thousand requests, it periodically resets the counters to detect distributional drift. When the average SLO attainment falls below a threshold and observed hit rates diverge from their expected values, an update cycle is triggered: re-profiling query access patterns, rerunning the latency-bounded partitioning algorithm, generating shards, and loading the updated indices onto GPUs.
All stages, from profiling to loading, complete in under a minute, allowing updates to run in the background. At the per-shard level, index generation and loading take less than ten seconds. The detailed timing breakdown for each stage is shown in Figure 9. While a GPU shard is being refreshed, the router temporarily redirects queries for those clusters to CPU paths, preserving the service continuity. Once the updated shard is loaded, routing automatically returns to the GPU.
Per-cluster updates are avoided because clusters are stored contiguously to enable high-bandwidth access. Since clusters vary in size, updating clusters individually would lead to memory fragmentation and inefficient data placement. Instead, VectorLiteRAG performs full-shard updates, as migration of each shard takes only a few seconds, providing robustness and simplicity.
According to our observations, profiling with only 0.5% of the queries from a separate training set successfully captured the distribution of 10M ORCAS queries. We therefore assume that a single index update can sustain stable service for roughly one hour under steady traffic, given the system throughput measured in our experiments.
Figure 9: Time consumed for re-building the GPU index shards using updated query access data. Numbers above the bars denote the search time SLO constraints applied for the system.
V Methodology
V-A Experiment Setup
@@
Datasets and Models. We use two datasets: Wiki-All and ORCAS. We construct the IVF index following the configuration guidelines provided by the Faiss library. The Wiki-All [37] vector database contains 88M 768-dimensional vectors derived from Wikitext [28] and Cohere Wikipedia embeddings, yielding a compressed IVF index with a footprint of 18GB. We also construct two additional indexes from chunked Wikipedia documents [40] using the Stella [42] embedding model of dimensions 1024 and 2048, and queries from the Microsoft ORCAS dataset [5]. ORCAS consists of real Bing queries and preserves duplicates to reflect realistic query distributions. The ORCAS 1K and ORCAS 2K indexes occupy 40GB and 80GB of memory, respectively.
Our retrieval pipeline builds on Faiss v1.9.0 [6, 17], with internal extensions for flexible nprobe settings and dispatcher callbacks. The overall system, including the profiler and latency-aware scheduler, is implemented in Python.
For generation, we evaluate three models—Llama3-8B, Qwen3-32B, and Llama3-70B [9, 41]—served using vLLM v0.9.1 [18]. The retriever and LLM run as separate subprocesses, with the main process coordinating request generation and document fetching to integrate the full RAG pipeline.
To evaluate system performance, we sample queries from a dedicated test set that is disjoint from the profiling set. The request arrival process follows a Poisson distribution, a commonly adopted modeling choice in prior work [18, 45, 31]. For each query, the top-25 documents are retrieved, and a 1024-token input is constructed and passed to the LLM, which then generates a 256-token output, following the setup in [35]. The initial nprobe is set to 2048, which is sufficient to achieve an average retrieval quality of 0.91 Normalized Discounted Cumulative Gain (NDCG) [39] at 50.
@@
TABLE I: SLO target values used in the main evaluation
@@
System Configuration. We conduct our experiments on two types of nodes, each equipped with eight NVIDIA GPUs. The L40S node includes L40S GPUs with 48GB GDDR memory and dual Xeon 6426Y CPUs. The H100 node uses H100 GPUs with 80GB HBM and Xeon Platinum 8462Y CPUs. We use the L40S node for smaller models (Llama3-8B), while larger models requiring model parallelism (Qwen3-32B, Llama3-70B) are run on the H100 node for maximum throughput.
@@
Baseline Configurations. We compare VectorLiteRAG against several key baselines. Since VectorLiteRAG builds on FAISS, we use vanilla FAISS-CPU IVF FastScan (CPU-Only), FAISS-GPU IVF on a dedicated GPU (DED-GPU), and a sharded FAISS-GPU IVF index distributed across all GPUs (ALL-GPU). To further demonstrate the strength of our approach, we also compare against HedraRAG [11] in section VI-D, which also uses a skew-aware caching strategy.
VI Evaluations
VI-A Performance Model and Hit Rate Estimator